### v1

# Are we there yet? A brief survey of Music Emotion Prediction Datasets, Models and Outstanding Challenges

Abstract
Deep learning models for music have advanced drastically in the last few years. But how good are machine learning models at capturing emotion these days and what challenges are researchers facing? In this paper, we provide a comprehensive overview of the available music-emotion datasets and discuss evaluation standards as well as competitions in the field. We also provide a brief overview of various types of music emotion prediction models that have been built over the years, offering insights into the diverse approaches within the field. Through this examination, we highlight the challenges that persist in accurately capturing emotion in music. Recognizing the dynamic nature of this field, we have complemented our findings with an accompanying GitHub repository. This repository contains a comprehensive list of music emotion datasets and recent predictive models [^1] https://github.com/AMAAI-Lab/awesome-MER/.
1 Introduction
Music has long been revered for its profound ability to evoke and convey emotions, transcending cultural, linguistic, and geographical barriers. Researchers from various fields have been captivated by the intricate interplay between music and human emotions for decades. Classic works such as Meyer’s paper [1] and pioneering studies conducted by scholars such as Seashore [2] and Hevner [3] in the early 20th century laid the groundwork for understanding the emotional impact of music. However, understanding and quantifying this poses a significant challenge due to its multifaceted nature [4]. In this paper, we provide an overview of the current state-of-the-art along with challenges and directions for future work.
With the advent of technology and data-driven methodologies, new avenues have opened up for exploring the complex relationship between music and emotions, such as deep learning models. Despite these advancements, accurate predictions of emotions from music remains elusive. This is due to a number of reasons, including the subjective, personal nature of emotional perception, as well as bias and limitations in the current datasets, and challenges in benchmarking. We discuss these and other challenges at length in Section 5.
Despite these challenges, the potential applications of Music Emotion Recognition (MER) systems are vast and varied. In the healthcare domain, we may find personalized music recommendation systems that guide a listener to different emotional states [5], or we may even see MER systems used to build large-scale datasets on which we can train generative music AI systems that can be controlled to generate music with specific emotions [6, 7]. In the same line of thought, [8] developed a Brain-Computer Interface system that can provide musical feedback about the listener’s current emotional state and subsequently influence this state. Additionally, MER systems may help facilitate emotional analysis during music composition, interactive experiences in media and entertainment, as well as inform market research [9]. The implications of understanding music and emotions extend across diverse domains. In general, following the idea of positive psychology [10], understanding music emotions can be used to enhance the user experience when designing various systems.
In this paper, we do not aim to provide a comprehensive overview of MER models, instead, we focus on discussing datasets, evaluation approaches, and identifying key challenges as well as future directions. We only briefly touch upon some of the more recent MER models. For a more comprehensive overview of MER machine learning models, the reader is referred to [11, 12, 13, 9, 14].
In the next section, we provide an extensive overview of the available emotion-annotated music datasets. This is followed by a discussion of the evaluation practices in the field (Section 3). After that, we describe a selected list of recent models and approaches in Section 4. Finally, Section 5 dives into the remaining challenges and future direction for the field of MER, followed by a general conclusion.
Table 1: Overview of Music Emotion Datasets (Sorted by Year).
Static/ Perceived/
Dataset Year # of instances Length Type Categorical Dimensional Dynamic Induced
MoodsMIREX [15] 2007 269 30s MP3 5 labels - Static Perceived
CAL500 [16] 2007 500 full MP3 174 labels - Static Perceived
Yang-Dim [17] 2008 195 25s WAV - Russell Static Perceived
MoodSwings [18] 2008 240 15s MP3 - Russell Dynamic Perceived
NTWICM [19] 2010 2,648 full MP3 - Russell Static Perceived
Soundtracks [20] 2011 470 15s-1m MP3 6 labels 3 dimemsions Static Perceived
DEAP [21] 2012 120 60s YouTube id - Russell Static Induced
Panda et al.’s dataset [22] 2013 903 30s MP3, MIDI 21 labels - Static Perceived
Solymani et al.’s dataset [23] 2013 1000 45s MP3 - Russell Both Perceived
Emotify [24] 2016 400 60s MP3 GEMS - Static Induced
Moodo [25] 2016 200 15s WAV - Russell Static Perceived
CH818 [26] 2017 818 30s MP3 - Russell Static Perceived
4Q-emotion [27] 2018 900 30s MP3 Quadrants - Static Perceived
MediaEval DEAM [28] 2018 2,058 45s MP3 - Russell Both Perceived
PMEmo [29] 2018 794 full MP3 - Russell Both Induced
RAVDESS [30] 2018 1,012 full MP3, MP4 5 labels - Static Perceived
DMDD [31] 2018 18,644 full Audio, Lyrics - Russell Static Perceived
MTG-Jamendo [32] 2019 18,486 full MP3 56 labels - Static Perceived
VGMIDI [33] 2019 200 full MIDI - Russell Dynamic Perceived
Turkish Music Emotion [34] 2019 400 30s MP3 4 labels - Static Perceived
EMOPIA [35] 2021 1,087 30s-40s Audio, MIDI Quadrants - Static Perceived
MER500 [36] 2020 494 10s WAV 5 labels - Static Perceived
Music4all [37] 2020 109,269 30s WAV - 3 dimensions Static Perceived
CCMED-WCMED [38] 2020 800 8-20s WAV - Russell Static Perceived
MuSe [39] 2021 90,001 full Audio - Russell (V-A-D) Static Perceived
HKU956 [40] 2022 956 full MP3 - Russell Static Induced
MERP [41] 2022 54 full WAV - Russell Both Perceived
MuVi [42] 2022 81 full YouTube id GEMS Russell Both Perceived
YM2413-MDB [43] 2022 699 full WAV, MIDI 35 labels - Static Perceived
MusAV [44] 2022 2,092 full WAV Russell Static Perceived
EmoMV [45] 2023 5,986 30s WAV 6 labels - Static Perceived
Indonesian Song [46] 2023 476 full WAV 3 labels - Static Perceived
TROMPA-MER [47] 2023 1,161 30s WAV 11 labels - Static Perceived
EMMA [48] 2024 364 30s-60s WAV GEMS - Static Perceived
SiTunes [49] 2024 300 full WAV - Russell Static Induced
2 Datasets
Table 1 presents an extensive overview of emotion-annotated music datasets. These datasets vary in size, annotation granularity, and focus on either perceived or induced emotions. Before delving deeper into the datasets, it is essential to understand the emotion representations commonly referenced in the field.
Emotion representations One of the earliest emotion representation models in music is Hevner’s affective ring [50], developed in 1936. Based on extensive experimental studies, Hevner’s model categorizes music emotions into eight fundamental categories: dignified, sad, dreamy, serene, graceful, happy, exciting, and vigorous. In current-day MER research, we see that Russell’s Circumplex Model of Affect [51] is widely used. This model characterizes emotions along two dimensions: valence and arousal. Valence represents the degree of positive or negative emotion, while arousal reflects the intensity of emotion, ranging from passive to activated states. Russell’s original model contained a third dimension: dominance [51]. This dimension is typically omitted as it can be hard to annotate, although some researchers have argued to re-include it [52]. Russell’s model is a dimensional model, as it consists of continuous values along multiple dimensions (valence/arousal). Hevner’s model, on the other hand, is categorical as it consists of discrete emotion labels.
Thayer’s two-dimensional model [53] focuses on energetic arousal and tense arousal as the primary dimensions of emotion. Thayer suggests that valence can be inferred from the combination of energetic and tense arousal levels.
Categorical models include the Geneva Emotional Music Scales (GEMS) [24]. This model was specifically designed for music-induced emotions and consists of 45 emotion tags grouped into nine categories, including amazement, solemnity, tenderness, nostalgia, calmness, power, joyful activation, tension, and sadness. The datasets listed in Table 1 use a variety of emotion representation models, with Russell’s model being the most popular dimensional representation. Some datasets do not use a specific emotion model, for instance, the MTG-Jamendo dataset consists of freely assigned tags by the listeners, resulting in a diverse and comprehensive set of 56 tags, spanning from ‘melancholic’ to ‘upbeat’. A subset of this dataset is used for the ‘Emotion and Theme Recognition in Music’ Task of MediaEval. In Table 1, the number of tags used to represent the emotions are listed in the column ‘Categorical’.
Static versus dynamic Regardless of which emotion representation model is being used, we notice two fundamentally different approaches: static versus dynamic annotations. In a static setting, the listeners indicate the emotion for the entire song or fragment. In a dynamic annotation setting, the listener continuously indicates the emotion throughout the song or fragment. For instance, the MTG-Jamendo dataset [32] offers categorical tags for each full-length song. The annotation is static, but multiple tags are allowed per song. The MoodSwings dataset [18], on the other hand, offers dynamic annotations of valence/arousal for every second of the musical fragments. Finally, some datasets offer both, for instance, MERP [41] provides both static GEM labels for the entire song, as well as dynamic valence/arousal ratings for every 1s of a song.
Induced versus perceived Emotion labels in the datasets can represent perceived or induced emotion. Perceived emotions are emotions that listeners consciously recognize while listening to the music. Induced emotions, on the other hand, are emotions that listeners experience as a result of listening to the music, but are not necessarily consciously recognized. Most of the datasets in Table 1 use perceived emotion labels, which are easier to annotate. For induced emotion labels, researchers have used psycho-physiological measurements ranging from electromyogram (EMG), volume pulse (BVP), electrocardiograms (ECG), skin conductance, respiration rate, heart rate, to electroencephalograms (EEG). There have been many studies in psychology that use such biosensors to explore how music can influence our emotions [54, 55, 56, 57, 58]. Since these studies include medical data, the datasets are not often public. However, there are a few datasets with music and its induced emotions. Firstly the DEAP dataset [21] includes EEG, facial video recordings, as well as peripheral physiological signals that were recorded while they watched music videos. In addition, they collected perceived emotion ratings in terms of arousal, valence, like/dislike, dominance and familiarity. Second, the HKU956 dataset [40] records five kinds of physiological signals (i.e., heart rate, electrodermal activity, blood volume pulse, inter-beat interval, and skin temperature) of participants as they listen to music, along with reported emotions in the arousal and valence dimensions. Finally, SiTunes [49] includes physiological signals (i.e., heart rate, activity intensity, activity step, and activity type) measured both before and after music listening, alongside environmental data (i.e., time of day, weather information, and location) recorded during users’ daily lives. The MER models that can predict induced emotions have various medical applications, such as curating playlists to guide patients to different emotional states [5]. An interesting study by [59] concludes that music tends to induce the emotion that is perceived, enabling researchers to use perceived emotion datasets for developing emotion-inducing models.
Modalities Music comes in many formats, the most common one being ‘audio’. Audio files can either be raw waveforms or compressed .mp3 files. However, we should not neglect the MIDI format, a popular format still often used by music producers, composers and performers. Whereas most datasets contain audio files, only a handful focus on MIDI: 1) the VGMIDI dataset [33], which contains continuous valence/arousal ratings for MIDI files of piano arrangements for video game soundtracks; 2) the Panda et al.’s dataset [22] (used for the MIREX competition), which provides a diverse collection of audio clips, lyrics, and aligned MIDI files, contains labels for five emotion clusters derived from a cluster analysis of online tags; and 3) the EMOPIA dataset [35], which contains paired piano music audio with MIDI that has emotion annotations of the four high/low valence/arousal quadrants.
Many sources can elicit emotion. For instance, video, or lyrics may also affect our perceived or induced emotions. Some datasets offer alternative multimedia streams such as emotion-rated music videos from a variety of genres, including pop, rock, classical, and jazz as in the DEAP dataset [21]; or text of lyrics with the annotated music in the DMDD dataset [31]. Finally, the MuVi dataset [42] even offers isolated modality ratings for music videos. In this study, the raters were presented with either the music video, the music alone, or the muted video. The final dataset contains ratings for each of these modalities separately as well as together. This allows [42] to build a model on pure isolated modalities, which proved to be more accurate than a traditional model.
Dataset size Compared to affective datasets available in other domains (e.g., the Sentiment140 dataset [60] in NLP, which contains 1.6 million tweets annotated as positive or negative), the size of the available datasets is still very limited, with the largest dataset containing 109,269 instances. In addition to the number of instances in datasets, we also notice a difference in the length of the instances. A number of datasets (e.g. MERP [41], VGMIDI [33], and HKU956 [40]) offer ratings for full-length songs. In datasets with dynamic ratings throughout the song, this may provide a means for researchers to analyse how our emotions evolve throughout a song. Other datasets focus on short fragments, often ranging from 30 seconds to 1-minute fragments (e.g. DEAM [28], EMOPIA [35], and EMMA [48]), but even as short as 10s as is the case for the MER500 dataset [36].
In sum, there is a variety of emotion-annotated datasets available as shown in Table 1. They differ in terms of emotion representation models, as well as music format and perceived versus induced emotion labels. The challenges and future research directions related to datasets are discussed in detail in Section 5.
3 Evaluation Protocols
Evaluation metrics play a crucial role in assessing the performance of MER systems. Depending on which type of emotion ratings are being predicted (dimensional versus categorical), the evaluation metrics change as we are dealing with a regression or a classification task respectively. Commonly used evaluation metrics for categorical MER systems include accuracy, precision, Area under the Curve (AUC), and confusion matrices. In the case of regression MER models, we see metrics such as precision, recall, F1-score, mean squared error (MSE), $R^{2}$, and Pearson correlation coefficient [61]. These metrics provide insights into the effectiveness of MER models when it comes to emotions represented by dimensional models.
Evaluating MER systems goes beyond just defining a common metric. Due to the inherent differences between datasets, a comparison across datasets is often not possible. For an in-depth discussion on this, the reader is referred to Section 5. Within one dataset, however, it is possible to establish benchmarks and compare the performance of different models, provided that the train/test split is shared. There are some initiatives to facilitate the comparison between models, such as competitions, as well as individual papers that offer clear data splits and metrics [62, 63, 64, 65, 66].
Competitions and challenges provide valuable platforms for evaluating and comparing the performance of MER systems. These initiatives often involve standardized datasets, evaluation protocols, and metrics, enabling researchers to benchmark their algorithms against state-of-the-art methods. Existing benchmarking initiatives and competitions in MER include the ‘Audio K-POP Mood Classification’ task in MIREX (Music Information Retrieval Evaluation eXchange) (last organized in 2019) [^2] https://www.music-ir.org/mirex/wiki/2019:Audio_K-POP_Mood_Classification/ , the ‘Emotion in Music’ task in MediaEval (last organized in 2015) [^3] http://www.multimediaeval.org/mediaeval2015/emotioninmusic2015/, and the ‘Emotion and Theme Recognition in Music Using Jamendo’ in MediaEval (last organized in 2021) [^4] https://multimediaeval.github.io/2021-Emotion-and-Theme-Recognition-in-Music-Task/.
Given the noisiness of labels in music-emotion datasets [41], one could also argue to include listening tests to verify the accuracy of the predicted labels, or even to perform extrinsic evaluation through a downstream task. The latter could be implementing MER in an emotion-inducing system, e.g. by performing playlist recommendations based on predicted emotions from music in the dataset, and subsequently evaluating if the listeners’ emotion indeed reaches the target emotion.
4 Models and approaches
In this section, we briefly touch upon some existing MER models. This is by no means a complete overview, for this, the reader is referred to other survey papers [11, 12, 13, 9, 14]. The aim of this section is merely to point out the current state-of-the-art and various approaches over the last few years.
Some of the earliest attempts at music emotion prediction involved rule-based approaches and hierarchical frameworks. For instance, Feng et al. [67] used Computational Media Aesthetics (CMA) to analyze tempo and articulation, mapping them into four mood categories: happiness, anger, sadness, and fear. They achieved a total precision of 67% and a total recall of 66%. Lu et al. [68] developed a hierarchical framework for automatically extracting music emotion from acoustic data. They employed music intensity to represent the energy dimension of Thayer’s model while using timbre and rhythm to capture the stress dimension. They achieved an average accuracy of mood detection of up to 86.3%. These early models laid the groundwork for subsequent advancements in MER, paving the way for the adoption of more sophisticated techniques, including deep learning approaches.
In the last few years, various deep-learning MER models have been developed. Many of these leverage common techniques for time-series data, including Recursive Neural Networks (RNNs) such as Long-Short Time Memory networks (LSTMs) [69, 42], and more recently, Transformer architectures [70, 46]. Trying to identify thé current state-of-the-art MER model is a near-impossible task, due to reasons discussed in the previous section, for instance, the difference in labels across datasets. Looking at a recent competition, the 2021 Emotion and Theme Recognition in Music Using MTG-Jamendo dataset, we see that the best-performing model [66] reaches an PR-AUC-macro of 0.150872. This model uses a Convolutional neural networks (CNN) architecture.
A distinction between the various models can be made based on the input modality. Some models are exclusively based on MIDI such as the multi-task architecture proposed by [71], and as such use a token-based representation. Given the limited size of MIDI datasets, the accuracy of such models is limited, e.g., the model by [71] reaches 67.56% accuracy when predicting between four emotion classes. Most of the existing MER models are based on audio, and hence they take as input raw audio. This is often converted into Mel-spectrograms which are then processed through convolutional neural networks (CNNs) [72], or the audio could be directly processed through a WaveNet architecture, which is a type of temporal CNN [73]. In recent years, audio embeddings pretrained on large-scale datasets have become available, providing researchers with tools to enable transfer learning. For instance, [74] incorporated a VGGish network with Transformers, achieving mean squared error (MSE) and Pearson correlation coefficient (PCC) values of 0.117 and 0.655 for arousal, respectively, and 0.170 and 0.575 for valence, respectively.
Alternatives to these approaches include directly extracting spectral features (e.g., MFCCs, spectral centroids) with libraries such as OpenSmile, which has a configuration file specifically for the emotion recognition task [75]. [62] uses this approach and achieves $R^{2}$ scores of 0.378 and 0.638 for predicting dynamic valence and arousal values, respectively, on the MediaEval dataset.
Finally, musically meaningful features may be included, such as Rhythmic features (e.g., tempo, beat histogram), or note features (e.g., pitch), as implemented by Shi et al. [76] and Panda et al. [27], respectively. Shi et al. achieved a precision of 92.8% for 4 emotion categories (calm, sad, pleasant, and excited), while Panda et al. achieved an F1-score of 76.0% for 4 emotion categories (Quadrants).
Other input modalities may include text (lyrics), or video. In the case of the former, some models extract the sentiment from the lyrics using Natural Language Processing (NLP) tools [77], or they use the entire lyrics with an embedding model [70]. These features are then combined with audio-based features or analyzed independently to predict the emotional content of music. Including the lyrics does not always improve the sentiment prediction, particularly in predicting arousal [77], however, [31] did manage to increase the performance of a MER model by using a word2vec [78] embedding trained on 1.6 million lyrics. Finally, models that include video modalities typically use pretrained networks to capture image and video features, and thus improve model performance. For instance, [45] uses ResNet-50 [79] and FlowNetS [80], respectively.
We have only provided a glimpse into the existing MER models in this section. From the performance of the various models, we see, however, that much improvement can still be made. We discuss some of the remaining challenges in the next section.
5 Challenges and future directions
Whereas the first publications on music and emotion surfaced in the 1930s [50], we have not yet reached the stage where models can reach human-like performance (e.g. $>$ 95% accuracy). The field of MER still faces several challenges and various opportunities for future exploration remain, ranging from overcoming data limitations to the integration of emerging technologies. Understanding and addressing these challenges is crucial for advancing the field and unlocking its full potential.
Dataset limitations One of the primary challenges in MER is the scarcity of large, diverse, copyright-cleared, emotion-annotated datasets. Limited datasets hinder the development and evaluation of robust MER models, leading to potential biases and generalization issues. We have recently seen advances in this area with the release of larger datasets such as MTG-Jamendo [32], Music4all [37], MuSe [39], which contain 18k, 109k, and 90k instances respectively. Whereas MTG-Jamendo and MuSe are available under a Creative Commons licence, however, Music4all contains copyrighted tracks. In addition, when exploring datasets, we notice that they are often skewed towards one particular genre. For instance, the DEAM dataset [28] consists mostly of rock and electronic music genres, while the VGMIDI dataset is focused solely on video game soundtracks.
To deal with the current dataset size limitations, techniques such as unsupervised pretrained may be helpful. With this technique, latent representations are first learned using unlabeled datasets. This evolution has led to some available large-scale audio encoders such as the Variational Auto Encoder used in [81], and the AST Audio Spectrogram Transformer presented in [82], as well as various recently developed neural audio encoders (e.g. Descript Audio Codec [83]). These novel pretrained representations may help deal with the limiting size of emotion-annotated datasets.
Subjective labels The subjective and variable nature of emotion perception also poses a significant challenge. Emotions are inherently complex and subjective, varying across individuals, cultures, and contexts [41]. For instance, researchers have observed significant dependence between the number of years of musical training [84, 41], gender [42], familiarity with songs [42], culture [85, 41], genre preference [41], and even age [41]. The latter study makes an argument for including profile information in the emotion prediction model to personalize the predictions. However, not many datasets include this type of information other than MERP.
In addition, many datasets have a cultural bias, meaning that they are annotated by people from the same culture, often simply because of the locality of the experiment or language constraints. However, it has been established that people from different countries or cultural backgrounds have different perceptions of emotion for the same music fragment [86]. For instance, [41] have raters from both the US as well as India, and see a significant difference in their annotations. HKU956 [40] go even further and include the rater’s responses to a personality test: ‘Ten Item Personality Measure’. In future work, one could develop datasets annotated by annotators from different cultural backgrounds, that include this information about the raters. The personalized nature of emotion ratings can cause a low inter-rater reliability, a metric of agreement between raters often calculated using Cronbach’s Alpha [87].
Noisy labels When creating datasets, an additional challenge arises: it can be hard to identify the emotion perceived from music, especially when working with valence and arousal models. In fact, Russel’s model [51] included a third dimension: dominance. This dimension is typically omitted because of the ambiguity in annotation. In general, categorical models, although less precise, are often easier to annotate [59]. This personal variability and ambiguity in emotion labels introduces noise in dataset labels, causing low inter-rater reliability, and making it challenging to train accurate and reliable MER models.
Machine learning models may also offer useful techniques to deal with noisy labels, as discussed in the survey by [88]. These could include Noisy Graph Cleaning (NGC) [89], Joint Training with Co-Regularization (JoCoR) [90], and Robust Curriculum Learning (RoCL) [91].
Annotation interfaces When creating a static dataset with static annotations, some standard tools can be used, including PsyToolkit [92]. However, to create any larger-scale dataset, the annotations are often done through crowdsourcing services such as Amazon Mechanical Turk [^5] https://www.mturk.com (e.g. for MERP [41], DEAM [28], and Solyman’s dataset [23]). These services offer access to a large ‘army’ of annotators, which may come at the cost of accuracy. There are, however, a number of techniques that can be used to filter out some of this annotation noise. This includes limiting the annotations to Master raters (raters with a known track record that often work at a premium price) [41], by including qualification tasks to assess participants’ understanding of the dimensional model [23], or by using multiple ground truth questions that are shown to all raters [41]. Finally, inter-rater reliability can be used to filter out low-quality annotations [41]. When doing this, it is important to keep in mind personal characteristics, which may cause different people to rate music differently. Hence, inter-rater reliability should ideally be calculated by taking into account the rater’s profile features.
An additional difficulty is that the amount of available interfaces for music emotion annotation is very limited. Especially when it comes to time-continuous annotations of valence and arousal. [41] released their dynamic annotation interface [^6] https://github.com/dorienh/MERP that hooks into Amazon mTurk. Kim et al. [18] also introduced, an annotation interface called MoodSwings, designed to record dynamic emotion labels.
Benchmarking The ImageNet competition has been instrumental in establishing a clear performance benchmark among computer vision models [93]. While there have been similar competitions for music emotion prediction (see Section 3), none of these competitions ran in the last three years, indicating the lack of a current benchmark for MER systems. To establish benchmarks on individual datasets, we have to revert to individual model papers, such as [28, 23]. It is unfortunately not always clear which train/test split these systems use, making direct comparisons hard. One way to facilitate an easy overview would be to leverage Leaderboard features on popular websites such as Papers With Code [^7] https://paperswithcode.com/sota and HuggingFace [^8] https://huggingface.co/docs/competitions/en/leaderboard.
Finally, it is also hard to make comparisons across datasets, as many of them use different emotion representations. While some of these differences may be bridged by using synonyms, it is currently not a common practice to compare across datasets. However, if we are able to merge different emotion representations and train across multiple datasets, then we can start leveraging larger-scale models. Bridging between datasets that use continuous representations and categorical models is the least evident. [94] have identified a mapping between arousal/valence and categorical labels. This has been used by [6] to facilitate emotion-controlled lead sheet generation, and it may offer a way to merge continuous and categorical datasets into a larger-scale dataset for music emotion.
MIDI The forgotten format in MER. Emotion originates from many different aspects of the music, including the tonal tension [7], instrumentation and timbre [95], production quality [96], expressiveness of the performance [97], harmony [98]. The symbolic MIDI format only captures parts of these [12], which may explain the lack of models for predicting emotion from MIDI. Another reason may simply be the lack of emotion-annotated datasets (three in total). Even though MIDI is an abstraction of music, it is still a widely used format by music producers and performers and warrants its own models for emotion prediction. Even more, such datasets may enable generative music systems (which are typically trained on MIDI) to be controlled by emotion [6, 99].
Multimodal predictions Our sensory input is multimodal. As such, some of the emotion-annotated datasets enable us to look at multiple modalities. For instance, DEAP [21] offers biofeedback data, i.e. EEG recordings, and frontal face videos from participants. Similarly, HKU956 [40] offers physiological signals including heart rate, electrodermal activity, blood volume pulse, inter-beat interval, and skin temperature. SiTunes [49], on the other hand, provides physiological and environmental situation recordings collected via smart wristband devices.
These biological data can serve as the induced emotion labels, e.g. EEG signals can be translated into human emotions [100]. Increased datasets with different types of physiological signals enable researchers to focus on creating models for induced versus perceived emotion detection. This offers new avenues to use biofeedback in music emotion mediation applications through smart devices.
Sometimes other emotion-inducing modalities are present, such as video, or lyrics. In this case, it is important to consider the influence of each of these modalities. In the case of video and music, Chua et al. [42] have studied the influence of exposing participants not only to the music but also the muted video, as well as the music videos. They found that the music modality explains most of the variance in arousal values, and both music and video modalities explain the variance in valence values. Phuong et al. [74] explore the influence of using only audio features and only video features to predict emotions from movie fragments. They found that the prediction is most accurate when both modalities are used. However, when predicting from a single modality, the audio model is most accurate.
Real-time Many of the currently existing models are not implemented as an easy-to-use library, nor are they quick to run. They often require a GPU and may take several minutes to run. There are use cases, however, for real-time emotion recognition systems, as they would be able to integrate into therapeutic emotion detection systems [5], mood guidance playlist systems, as well as more commercial systems such as advertisement targeting systems.
In sum, the challenges mentioned above provide direct opportunities and future directions to further advance the exciting field of MER.
6 Conclusion
MER is a promising field with various practical applications. With the rise of large-language models, we have seen impressive performance in various tasks. The field of MER, however, still seems to be lagging. Given the importance of large training datasets to facilitate the training of LLMs, we provide a comprehensive overview and discussion of the existing datasets for MER.
We also explore current state-of-the-art models and dive into evaluation methods such as metrics as well as competitions, leaderboards, and benchmarks within the MER field. With this knowledge, we discuss the current challenges of the MER field at length and provide concrete future directions and emerging trends such as real-time systems and multimodal prediction systems.
In closing, this survey serves as a valuable resource for the MER community, offering insights into the current state-of-the-art, as well as a discussion of challenges and inspiration for future directions.
7 Acknowledgements
This work has received SEED funding from SUTD TL under grant number RTDS S 22 14 04 1.